If 6 science fair exhibits are chosen at random from 4 biology, 7 chemistry, and 9 physics exhibits to

go to the county finals, what is the probability that there will be 2 from each subject area?


1
Expert's answer
2020-11-10T19:45:00-0500

Biology = 4

Chemistry = 7

Physics = 9

Total = 20


Number of ways of choosing 6 science fair exhibit out of 20 = 20C6 = 38760


Number of ways of choosing 2 biology = 4C2 = 6


Number of ways of choosing 2 chemistry = 7C2 = 21


Number of ways of choosing 2 physics = 9C2 = 36


Total number of ways of choosing 2 each of Biology, chemistry and physics = 6 × 21 × 36 = 4536


Probability of choosing 2 each of Biology, chemistry and physics = "\frac{4536}{38760}" = 0.1170